# CLARIFID: Improving Radiology Report Generation by Reinforcing Clinically Accurate Impressions and Enforcing Detailed Findings

### v1

@@ Abstract
Automatic generation of radiology reports has the potential to alleviate radiologists’ significant workload, yet current methods struggle to deliver clinically reliable conclusions. In particular, most prior approaches focus on producing fluent text without effectively ensuring the factual correctness of the reports and often rely on single-view images, limiting diagnostic comprehensiveness. We propose CLARIFID, a novel framework that directly optimizes diagnostic correctness by mirroring the two-step workflow of experts. Specifically, CLARIFID (1) learns the logical flow from Findings to Impression through section-aware pretraining, (2) is fine-tuned with Proximal Policy Optimization in which the CheXbert F1 score of the Impression section serves as the reward, (3) enforces reasoning-aware decoding that completes “Findings” before synthesizing the “Impression” , and (4) fuses multiple chest X-ray views via a vision-transformer-based multi-view encoder. During inference, we apply a reasoning-aware next-token forcing strategy followed by report-level re-ranking, ensuring that the model first produces a comprehensive “Findings” section before synthesizing the “Impression” and thereby preserving coherent clinical reasoning. Experimental results on the MIMIC-CXR dataset demonstrate that our method achieves superior clinical efficacy and outperforms existing baselines on both standard NLG metrics and clinically aware scores.
Introduction
@@
Although recent studies have advanced both fluency and factual grounding in automated reports, several critical limitations persist: (1) heavy reliance on specialized modules or large annotated datasets that limit real-world applicability; (2) optimization for text-similarity metrics (e.g., BLEU, ROUGE) rather than clinical correctness; and (3) a predominant focus on the “Findings” section while neglecting the critical “Impression” section that directly drives diagnostic decision-making.
@@
Figure 1: Radiologist and CLARIFID follow the same clinical workflow—recording Findings first and then deriving the Impression—demonstrating the model’s alignment with expert practice.
@@
The most widely used datasets—MIMIC-CXR (Johnson et al. 2019) and IU X-ray (Demner-Fushman et al. 2016)—both provide reports with clearly separated Findings and Impression sections. Notably, in MIMIC-CXR, 82.4% of diagnostic labels are derived from the Impression section, highlighting its fundamental clinical importance. Yet most prior work in radiology report generation has prioritized generating detailed Findings with less attention to the diagnostic synthesis required for the Impression, creating a significant gap between research and clinical practice.
@@
We reframe radiology report generation as a structured clinical reasoning task that mirrors the radiologist’s workflow: first documenting detailed observations (“Findings”), then synthesizing them into diagnostic conclusions (“Impression”) (see Figure 1). The Impression serves as the “answer” that directly informs clinical decision-making, while Findings provide supporting evidence and context. Analogous to recent advances in language model reasoning—which show that optimizing for final-answer correctness can improve intermediate reasoning steps (DeepSeek-AI 2025)—we hypothesize that optimizing for clinical accuracy of the Impression section will encourage higher-quality Findings documentation through implicit learning of the relationship between observations and diagnoses.
@@
To further support high-quality report generation during inference, we introduce a controlled decoding strategy that prevents premature transition to the Impression section. By enforcing a minimum generation length for the Findings section through “next-token forcing”, we ensure that the model provides sufficient observational detail before synthesizing conclusions. This reasoning-aware decoding complements our training strategy, yielding a consistent clinical reasoning structure that enhances report interpretability and fidelity.
Furthermore, radiologists routinely inspect multiple CXR projections of the same patient to reach an accurate diagnosis, yet most existing RRG models still process a single image, restricting cross‑view diagnostic insight. To address this limitation, we introduce a flexible multi‑view fusion module capable of integrating any number of CXR images.
@@
  • We propose an Impression-centric reinforcement learning framework using CheXbert F1 scores as rewards, directly optimizing for clinical accuracy.
  • We introduce a reasoning-aware decoding strategy that first ensures comprehensive “Findings” generation before “Impression” synthesis and then, through report-level re-ranking, simultaneously elevates both the “Findings” and “Impression” sections.
  • We adapt a multi-view fusion architecture with attention masking for variable-length X-ray inputs to enhance clinical efficacy.
Figure 2: Overview of the CLARIFID framework. (1) Section-aware pretraining flags each Findings sentence with <next> and the Impression with <impression>, teaching the report’s two-step structure. (2) After a supervised warm-up, the model is fine-tuned by PPO, using CheXbert-F1 on the Impression as reward. (3) At inference, <impression> is withheld until at least $k$ <next> tokens have been generated; the model then samples $N$ complete candidate reports and outputs the top-scoring report.
Related Works
Radiology Report Generation
Recent advances in radiology report generation (RRG) have explored a wide range of architectures and training strategies to enhance clinical accuracy and textual coherence. Memory-based architectures (Chen et al. 2022b, a) aim to preserve long-range dependencies between findings and impression sections. Knowledge graph integration methods (Zhang et al. 2020; Liu et al. 2021; Huang, Zhang, and Zhang 2023; Li et al. 2023) inject structured medical relationships to guide generation. Contrastive learning techniques (Wang et al. 2022; Li et al. 2023, 2024) reduce dataset bias by distinguishing abnormal findings from normal background patterns. Object detector-based models (Tanida et al. 2023; Zhang et al. 2024) improve explainability by aligning textual phrases with anatomical regions. Prompt-driven frameworks (Jin et al. 2024) enhance controllability of content and style using auxiliary diagnostic signals. More recent studies have incorporated large language models (LLMs), fine-tuned using visual adapters or domain-specific bootstrapping strategies (Liu et al. 2024; Wang et al. 2023c).
Beyond supervised learning, reinforcement learning (RL) has been introduced to align generation objectives with clinically meaningful outcomes. Early work such as R2GenRL (Qin and Song 2022) used BLEU-based rewards to encourage visual-textual coherence. Subsequent studies shifted toward optimizing clinical correctness via entity-level supervision. For instance, factENT (Miura et al. 2021) measures entity-level matches based on exact or entailment-based criteria, while RadGraph (Jain et al. 2021) introduces a graph-based evaluation framework. (Delbrouck et al. 2022) used RadGraph scores as a reward signal to enhance factual consistency. Multi-objective RL methods such as MPO (Xiao et al. 2024) further model radiologist preferences as weighted reward vectors. These works highlight the growing trend of aligning generative models with clinical knowledge and diagnostic metrics—a foundation for our own approach that integrates multi-image encoding with clinically informed RL objectives.
Test-Time Scaling Strategies
@@
Parallel methods generate multiple candidate outputs and select the best one using techniques such as majority voting or best-of-$N$ sampling with a reward model (Wang et al. 2023a; Brown et al. 2024; Irvine et al. 2023; Chen et al. 2024; Song et al. 2024). These approaches improve output quality through sampling, without altering model weights.
Sequential approaches, in contrast, iteratively refine a single reasoning path. Starting with an initial response, the model revisits and improves its output based on predefined criteria such as logical consistency, task-specific constraints, or self-generated feedback. Techniques like self-refine (Madaan et al. 2023), reflexion (Shinn et al. 2023), and decomposition-based self-correction (Ferraz et al. 2024; Li et al. 2022), as well as chain-of-thought prompting (Wei et al. 2023), exemplify this approach. These methods deepen reasoning without requiring multiple samples, offering a compute-efficient path to improved performance. Budget forcing (Muennighoff et al. 2025) and adaptive injection decoding (Jin et al. 2025) further illustrate sequential scaling. These methods regulate inference-time computation by either forcibly terminating generation or prolonging reasoning through repeated insertion of control tokens, particularly when the model attempts to conclude prematurely. Unlike majority voting, whose performance often saturates with more samples, token forcing strategies can benefit from increased reasoning depth.
@@
Method
@@
Our CLARIFID framework integrates multi-view image encoding, structured pretraining, reinforcement learning, and controlled decoding to generate clinically faithful radiology reports. Figure 2 illustrates the complete framework, consisting of three main components: (1) section-aware pretraining, (2) impression-guided reinforcement learning, and (3) next-token forcing with best-of-N impression selection. In this section, we elaborate on each component in detail.
Model Architecture
@@
Both the policy and value networks are based on the GPT-2 architecture (Radford et al. 2019). Visual information is supplied by a multi-image encoder that fuses all available X-ray views of a study. Conditioned on these fused features, the policy network generates reports, while the value network estimates expected rewards for reinforcement learning. The two networks use different depth configurations to balance performance and computational efficiency. For additional architecture details of the policy and value networks, see Appendix.
Multi-Image Encoder
Radiologists usually examine several X-ray views of the same patient to reach an accurate diagnosis, yet most RRG systems still rely on a single image, limiting diagnostic insight. We therefore present a flexible multi-view fusion module that can handle a variable number of X-ray images from different studies within a batch.
For a batch of $B$ studies where study $i$ contains $K_{i}$ images, we first flatten all images into a single processing batch:
=$\mathcal{I}=\{I_{i,j}\mid i\in[1,B],j\in[1,K_{i}]\}$ (1)
where $I_{i,j}\in\mathbb{R}^{C\times H\times W}$ represents the $j$-th image from study $i$ with $C$ channels and spatial dimensions $H\times W$. We track the count vector $\mathbf{c}=[K_{1},K_{2},\ldots,K_{B}]$ to maintain image-study associations.
All images in $\mathcal{I}$ are processed through a shared SwinV2 vision encoder $E$:
$\mathbf{F}=E(\mathcal{I})\in\mathbb{R}^{|\mathcal{I}|\times N\times D}$ (2)
where $N$ is the number of visual tokens per image and $D$ is the feature dimension. We then split $\mathbf{F}$ according to $\mathbf{c}$ to obtain study-specific features $\mathbf{F}_{i}\in\mathbb{R}^{K_{i}\times N\times D}$, which are flattened into study-level representations $\mathbf{S}_{i}\in\mathbb{R}^{K_{i}\cdot N\times D}$.
To handle the varying lengths across the batch, we create padded representations:
@@
$\widehat{\mathbf{S}}$ $=\text{Pad}([\mathbf{S}_{1},\mathbf{S}_{2},\ldots,\mathbf{S}_{B}])\in\mathbb{R}^{B\times M_{max}\times D}$ (3)
$\mathbf{A}$ $=\text{Pad}([\mathbf{1}_{K_{1}\cdot N},\mathbf{1}_{K_{2}\cdot N},\ldots,\mathbf{1}_{K_{B}\cdot N}])\in\{0,1\}^{B\times M_{max}}$ (4)
where $M_{max}=\max_{i}(K_{i}\cdot N)$ is the maximum number of visual tokens in the batch, and $\mathbf{1}_{K_{i}\cdot N}$ represents a vector of ones with length $K_{i}\cdot N$. The resulting matrix $\mathbf{A}$ serves as an attention mask, with $A_{i,j}=1$ for valid tokens and $A_{i,j}=0$ for padding. This pair $(\widehat{\mathbf{S}},\mathbf{A})$ is then passed to the decoder networks.
Section-Aware Pretraining
To explicitly represent the two-stage structure of radiology reports—detailed observations (Findings) followed by a diagnostic summary (Impression)—we introduce a section-aware pretraining scheme with special control tokens that flag the report’s logical structure.

### v4

@@ Abstract
Automatic generation of radiology reports has the potential to alleviate radiologists’ significant workload, yet current methods struggle to deliver clinically reliable conclusions. In particular, most prior approaches focus on producing fluent text without effectively ensuring the factual correctness of the reports and often rely on single-view images, limiting diagnostic comprehensiveness. We propose CLARIFID, a novel framework that directly optimizes diagnostic correctness by mirroring the two-step workflow of experts. Specifically, CLARIFID (1) learns the logical flow from Findings to Impression through section-aware pretraining, (2) is fine-tuned with Proximal Policy Optimization in which the CheXbert F1 score of the Impression section serves as the reward, (3) employs controlled decoding that completes “Findings” before synthesizing the “Impression”, and (4) fuses multiple chest X-ray views via a vision-transformer-based multi-view encoder. During inference, we apply a next-token forcing strategy followed by report-level re-ranking, ensuring that the model first produces a comprehensive “Findings” section before synthesizing the “Impression” and thereby preserving coherent clinical reasoning. Experimental results on the MIMIC-CXR dataset demonstrate that our method achieves superior clinical efficacy and outperforms existing baselines on clinical efficacy scores. The source code is available at https://github.com/IncheonYSH/CLARIFID.
1 Introduction
@@
Over the past few years, radiology report generation (RRG) has progressed along two axes that matter directly for clinical utility. The first of these concerns representation. Models strengthen long‐range clinical context and image–text alignment with memory or hierarchy enhanced encoder–decoders Chen et al. (2020, 2022). To curb hallucinations and bias toward frequent normalities, many systems inject structured priors: medical knowledge graphs, entity-relation schemas, or contrastive objectives that pull abnormal cues closer in representation space Zhang et al. (2020); Jain et al. (2021); Liu et al. (2021a); Huang et al. (2023); Li et al. (2023); Wang et al. (2022); Li et al. (2024). A complementary line explicitly grounds textual spans to anatomical evidence via detector-aligned attention or region-phrase alignment, improving localization and explainability Tanida et al. (2023); Zhang et al. (2024). Concurrently, LLM-based approaches leverage lightweight visual adapters and diagnosis-driven prompts to better capture radiology-specific style and higher-level reasoning without retraining massive backbones Jin et al. (2024); Wang et al. (2023c); Liu et al. (2024a). Multi-view CXR processing has also advanced; however, practical simplifications—such as fixing view configurations to rigid single/dual-view setups or reshaping reports to fit single-image paradigms—can inadvertently limit access to inter-view nuances and clinically informative comparison cues, leaving complementary projections under-leveraged in practice Liu et al. (2025); Wang et al. (2025).
The other axis concerns the training objective. Beyond cross-entropy on image–report pairs, reinforcement learning (RL) has been explored to align training with evaluation signals; however, rewards are often based on text-overlap metrics (BLEU/ROUGE/METEOR) that only partially capture clinical correctness Qin and Song (2022). To better reflect radiologists’ judgments, semantically informed objectives reward entity consistency or graph coherence—e.g. RadGraph based signals to reduce hallucinations and strengthen factual grounding Miura et al. (2021); Jain et al. (2021); Delbrouck et al. (2022). Recent work further considers multi-objective preference formulations to accommodate heterogeneous clinical criteria across conditions and report sections Xiao et al. (2024). In parallel, clinically grounded automatic labelers (e.g., CheXbert) have enabled scalable, label-space-aware assessment of clinical efficacy, motivating objectives that move beyond surface similarity toward decision-relevant accuracy Smit et al. (2020). Despite these advances, three gaps remain: reliance on supervision that requires labeling pipelines beyond standard image–report pairs, proxy-similarity optimization, and under-emphasis of the Impression—gaps that motivate our design.
@@
Figure 1: Radiologist and CLARIFID follow the same clinical workflow, recording the Findings first and then deriving the Impression. This demonstrates the model’s alignment with expert practice.
@@
The most widely used datasets, MIMIC-CXR Johnson et al. (2019) and IU X-ray Demner-Fushman et al. (2015), both provide reports in which the Findings and Impression sections are clearly separated. Notably, in MIMIC-CXR, 82.4% of diagnostic labels are derived from the Impression section Johnson et al. (2019), highlighting its fundamental clinical importance. Yet most prior work in radiology report generation has prioritized generating detailed Findings with less attention to the diagnostic synthesis required for the Impression, creating a significant gap between research and clinical practice.
@@
We reframe radiology report generation as a structured clinical reasoning task that mirrors the radiologist’s workflow: first documenting detailed observations (“Findings”), then synthesizing them into diagnostic conclusions (“Impression”) (see Figure 1). The Impression serves as the “answer” that directly informs clinical decision-making, while Findings provide supporting evidence and context. Analogous to recent advances in language model reasoning—which show that outcome-based optimization improves overall performance and can correlate with higher-quality intermediate steps Cobbe et al. (2021); Zelikman et al. (2024); DeepSeek-AI et al. (2025); Shen et al. (2025); Xie et al. (2025)—we hypothesize that optimizing for clinical accuracy of the Impression section will encourage higher-quality Findings documentation through implicit learning of the relationship between observations and diagnoses.
@@
To further support high-quality report generation during inference, we introduce a controlled decoding strategy that prevents premature transition to the Impression section. We delay “Impression” tokens with next‑token forcing until the “Findings” section reaches a preset length and then apply our section‑aware temperature schedule, using a higher temperature for Findings to surface detail and a lower one for Impression to maintain focus.
Furthermore, radiologists routinely review multiple projections from a single CXR study, but prior studies are often inaccessible Ge et al. (2012), and many existing RRG frameworks either overlook this reality Chen et al. (2020, 2022); Wang et al. (2023c); Jin et al. (2024) or adopt overly complex architectures Liu et al. (2025). To address this limitation, we introduce a flexible multi‑view fusion module capable of integrating any number of CXR images in one study.
@@
  • We propose an Impression-centric reinforcement learning framework, directly optimizing for clinical accuracy.
  • We introduce a reasoning‑aware decoding strategy specialized for radiology report generation that reflects both clinical intent and report structure.
  • We adapt a multi‑view fusion architecture to process variable‑length X‑ray inputs from a single study.
2 Related work
2.1 Reinforcement learning for RRG
We view reinforcement learning (RL) in radiology report generation (RRG) as a framework for optimizing non-differentiable, decision-oriented objectives that maximum-likelihood estimation (MLE) does not directly capture. A policy generates a candidate report, a learned or rule-based clinical reward evaluates its utility, and the policy is updated via policy-gradient or actor–critic methods. Preference-based alignment (e.g., RLHF/RLAIF/DPO) extends this mechanism by learning from human or AI preferences Ouyang et al. (2022); Lee et al. (2023); Rafailov et al. (2023); however, in clinical settings such feedback is costly, privacy-restricted, and often section-agnostic, making it difficult to enforce Impression-centric objectives.
Within RRG, early RL mainly optimized proxy text-overlap rewards (BLEU/ROUGE/METEOR) Qin and Song (2022), and subsequent work employed semantically grounded signals such as entity or graph consistency (e.g., RadGraph) to curb hallucinations and improve factual grounding Miura et al. (2021); Jain et al. (2021); Delbrouck et al. (2022). In contrast, we adopt a section-dependent objective that targets Impression-level clinical correctness via a CheXbert-derived signal Smit et al. (2020), while relying only on standard image–report pairs (no auxiliary labels or detectors). Optimization details are provided in Section 3.
2.2 Test-time scaling strategies
@@
Parallel methods generate multiple candidate outputs and select the best one using techniques such as majority voting or best-of-N sampling with a reward model Wang et al. (2023a); Brown et al. (2024); Irvine et al. (2023); Chen et al. (2024); Song et al. (2025). These approaches improve output quality through sampling, without altering model weights.
Sequential approaches, in contrast, iteratively refine a single reasoning path. Starting with an initial response, the model revisits and improves its output based on predefined criteria such as logical consistency, task-specific constraints, or self-generated feedback. Techniques like self-refine Madaan et al. (2023), reflexion Shinn et al. (2023), and decomposition-based self-correction Ferraz et al. (2024); Li et al. (2022), as well as chain-of-thought prompting Wei et al. (2022), exemplify this approach. These methods deepen reasoning without requiring multiple samples, offering a compute-efficient path to improved performance. Budget forcing Muennighoff et al. (2025) and adaptive injection decoding Jin et al. (2025) further illustrate sequential scaling. These methods regulate inference-time computation by either forcibly terminating generation or prolonging reasoning through repeated insertion of control tokens, particularly when the model attempts to conclude prematurely. Unlike majority voting, whose performance often saturates with more samples, token forcing strategies can benefit from increased reasoning depth.
@@
Figure 2: Overview of the CLARIFID framework. (1) Section-aware pretraining flags each Findings sentence with <next> and the Impression with <impression>, teaching the report’s two-step structure. (2) After a supervised warm-up, the model is fine-tuned by PPO, using CheXbert-F1 on the Impression as reward. (3) At inference, <impression> is withheld until at least $k$ <next> tokens have been generated; the model then samples NN complete candidate reports and outputs the top-scoring report. Refer to Figure 3 for details on the architectures of the policy and value models.
3 Method
@@
Our CLARIFID framework integrates multi-view image encoding, structured pretraining, reinforcement learning, and controlled decoding to generate clinically faithful radiology reports. Figure 2 illustrates the complete framework. In this section, we elaborate on each component in detail.
3.1 Model architecture
@@
Both the policy and value networks are based on the GPT-2 architecture Radford et al. (2019). Visual information is supplied by a SwinV2 Liu et al. (2021b)-based multi-image encoder that fuses all available X-ray views of a study. Conditioned on these fused features, the policy network generates reports, while the value network estimates expected rewards for reinforcement learning. The two networks use different depth configurations to balance performance and computational efficiency. A compact schematic with the module-level connections is provided in Figure 3.
Figure 3: Model architecture schematic diagram summarizing modules and data flow.
3.1.1 Multi-image encoder
Radiologists typically examine several X-ray views (e.g., PA, AP, LAT). To mirror this workflow, our model processes a variable number of images per study. For a batch of $B$ studies, where study $i$ has $K_{i}$ images, each image $I_{i,j}$ is encoded by a Swin Transformer encoder $E$. Let the last-stage feature map have spatial size $H_{f}\times W_{f}$ and channel $D_{\mathrm{img}}$. We apply a $1{\times}1$ convolution $W_{\mathrm{proj}}:\mathbb{R}^{D_{\mathrm{img}}}\!\to\!\mathbb{R}^{D}$ to match the model hidden size $D$, then flatten to $N{=}H_{f}W_{f}$ patch tokens, yielding $X_{i,j}\in\mathbb{R}^{N\times D}$.
For study $i$, we concatenate its $K_{i}$ image-token sequences $X^{\mathrm{concat}}_{i}=[X_{i,1};\,X_{i,2};\,\dots;\,X_{i,K_{i}}]\in\mathbb{R}^{(K_{i}N)\times D},$ so $M_{i}{=}K_{i}N$ tokens are available per study. Packing a batch and padding to $M_{\max}{=}\max_{i}M_{i}$ produces padded image feature $\widehat{\mathbf{S}}\in\mathbb{R}^{B\times M_{\max}\times D},$ and attention mask $\mathbf{A}\in\{0,1\}^{B\times M_{\max}},$ where $A_{i,m}{=}1$ denotes a valid visual token and 0 denotes padding. We feed $(\widehat{\mathbf{S}},\mathbf{A})$ directly to the text decoder’s cross-attention as keys/values and the encoder mask, respectively.
3.1.2 Policy network architecture
The policy network employs a decoder with 30 transformer layers, hidden size $D=768$, and 12 attention heads. It processes the padded image representations $\widehat{\mathbf{S}}\in\mathbb{R}^{B\times M_{\text{max}}\times D}$ and attention mask $\mathbf{A}\in\{0,1\}^{B\times M_{max}}$ through cross-attention at each transformer layer:
$h_{l}^{self}$ $=\text{SelfAttn}(h_{l-1})$ (1)
$h_{l}^{cross}$ $=\text{CrossAttn}(h_{l}^{self},\widehat{\mathbf{S}},\text{mask}=\mathbf{A})$ (2)
$h_{l}$ $=\text{FFN}(h_{l}^{cross})$ (3)
In the cross-attention operation, we use learnable projections $W_{Q},W_{K},W_{V},W_{O}\in\mathbb{R}^{D\times D}$. Queries, keys, and values are
@@ 3.1.2 Policy network architecture
$\mathcal{Q}\;=\;h_{l}^{\text{self}}W_{Q}\in\mathbb{R}^{B\times T\times D},\quad\mathcal{K}\;=\;\widehat{\mathbf{S}}W_{K}\in\mathbb{R}^{B\times M_{\max}\times D},\quad\mathcal{V}\;=\;\widehat{\mathbf{S}}W_{V}\in\mathbb{R}^{B\times M_{\max}\times D}.$
From the binary padding mask $\mathbf{A}$ we form a mask $\mathcal{M}\in\mathbb{R}^{B\times 1\times M_{\max}}$ with $\mathcal{M}_{i,1,m}{=}$-$\infty$ if $A_{i,m}{=}0$ and 0 otherwise (broadcast across $T$). The attention and output are
$\alpha\;=\;\mathrm{softmax}\!\left(\frac{\mathcal{Q}\mathcal{K}^{\top}}{\sqrt{D}}+\mathcal{M}\right)\in\mathbb{R}^{B\times T\times M_{\max}},\qquad\widetilde{H}_{l}\;=\;\alpha\mathcal{V}\in\mathbb{R}^{B\times T\times D},\qquad h_{l}^{\text{cross}}\;=\;\widetilde{H}_{l}W_{O}.$
The attention mask $\mathbf{A}$ ensures that for each study $i$, only the first $K_{i}\cdot N$ entries in row $i$ (representing valid visual tokens) influence the generation. This architecture enables clinically accurate reports conditioned on all available views, with the policy network outputting a distribution $P_{\theta}(y_{t}|y_{<t},\widehat{\mathbf{S}},\mathbf{A})$ at each step, where $h_{l}\in\mathbb{R}^{B\times T\times D}$(sequence length $T$) is the hidden state of layer $l$, $y_{t}$ is the target token at step $t$, $y_{<t}$ denotes the prefix tokens up to t1t{-}1, and $\theta$ collects all policy network parameters.
3.1.3 Value network architecture
The value network employs a decoder with 6 transformer layers, hidden size $D=768$, and 12 attention heads. It adopts the same cross-attention interface and inputs $(\widehat{\mathbf{S}},\mathbf{A})$ as the policy network, but replaces the language-modeling head with a position-wise linear value head. Let $h^{\text{final}}\in\mathbb{R}^{B\times T\times D}$ denote the final hidden states of the value decoder. We evaluate the value function $V_{\phi}$ at the autoregressive state $s_{i,t}=(y_{i,<t},\,\widehat{\mathbf{S}}_{i},\,\mathbf{A}_{i})$. The value head $f_{\phi}$ is an affine map applied independently at each token position with parameters $\mathbf{w}_{\phi}\in\mathbb{R}^{D}$ and $b_{\phi}\in\mathbb{R}$.
$\mathbf{v}=f_{\phi}(h^{\text{final}})=h^{\text{final}}\,\mathbf{w}_{\phi}+b_{\phi}\in\mathbb{R}^{B\times T},\qquad V_{\phi}(s_{t})=\mathbf{v}_{t}.$
Here $\phi$ collects all parameters of the value network.
3.2 Section-aware pretraining
We first warm up the image encoder and the lightweight aggregator on a 14-condition multi-label classification task. For a batch of $B$ studies with variable numbers of images, we reuse the padded visual tokens $\widehat{\mathbf{S}}\in\mathbb{R}^{B\times M_{\max}\times D}$ and the binary mask $\mathbf{A}\in\{0,1\}^{B\times M_{\max}}$. A lightweight Transformer encoder $f_{\mathrm{agg}}$ aggregates the tokens and a masked average pooling produces a study-level representation:
$\mathbf{H}=f_{\mathrm{agg}}(\widehat{\mathbf{S}},\,\mathbf{A})\in\mathbb{R}^{B\times M_{\max}\times D},\qquad\mathbf{h}_{i}=\frac{1}{Z_{i}}\sum_{m=1}^{M_{\max}}A_{i,m}\,\mathbf{H}_{i,m},\quad Z_{i}=\sum_{m=1}^{M_{\max}}A_{i,m}.$
Here, $m$ indexes the padded visual-token sequence (concatenated across all images), and $A_{i,m}{=}1$ if the $m$-th token of study $i$ is valid.
After masked average pooling, a linear classification head maps each $\mathbf{h}_{i}\in\mathbb{R}^{D}$ to 14 logits with parameters $W_{\mathrm{cls}}\in\mathbb{R}^{14\times D}$ and $\mathbf{b}_{\mathrm{cls}}\in\mathbb{R}^{14}$.
$\mathbf{z}_{i}=W_{\mathrm{cls}}\mathbf{h}_{i}+\mathbf{b}_{\mathrm{cls}}\in\mathbb{R}^{14},\qquad z_{i,c}=[\mathbf{z}_{i}]_{c}.$ (4)
We then convert CheXbert Smit et al. (2020) condition labels to binary targets, where positive is mapped to $1$ and blank, negative, and uncertain are mapped to 0. Let $\tilde{g}_{i,c}\in\{\texttt{blank},\texttt{positive},\texttt{negative},\texttt{uncertain}\}$ denote the original label. Define
$g_{i,c}=\mathbbm{1}[\tilde{g}_{i,c}=\texttt{positive}]=\begin{cases}1,&\tilde{g}_{i,c}=\texttt{positive},\\ 0,&\tilde{g}_{i,c}\in\{\texttt{blank},\texttt{negative},\texttt{uncertain}\}\end{cases}$ (5)
We optimize a per-condition weighted binary cross-entropy with weights $w_{c}$ estimated on the training split to mitigate class imbalance, where $w_{c}=\frac{n_{c}^{-}}{\,n_{c}^{+}}$ with $n_{c}^{+}$ and $n_{c}^{-}$ denoting the numbers of positive and negative instances for condition $c$ on the training split, and with $\sigma(\cdot)$ denoting the logistic function:
$\mathcal{L}_{\mathrm{cls}}=-\frac{1}{B}\sum_{i=1}^{B}\sum_{c=1}^{14}\Bigl[w_{c}\,g_{i,c}\,\log\sigma(z_{i,c})+(1-g_{i,c})\,\log\!\bigl(1-\sigma(z_{i,c})\bigr)\Bigr]$ (6)